please show step ny step solving for
100P-4000/100P=.3 for maintenance margin and stock price
Answer #1

100 P - 4000 = 30P

100P - 30P = 4000

70P = 4000

P = 4000 / 70

P = 57.14

The Price will be $ 57.14

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.